OpenStudy (anonymous):

Consider x^2 on an interval [0,1/2]. The Mean Value theorem suggest that there is a number c in (0.1/1/2) such that f'(c) is equal to a particular d.What is d?

OpenStudy (anonymous):

find the slope of the secant line that passes through (.1, .01) and (.5, .25)

OpenStudy (anonymous):

sorry, my browser crashed

OpenStudy (anonymous):

is the interval [0, 1/2] or [0.1, 1/2]?

OpenStudy (anonymous):

well, I'll assume [0, 1/2] then. Let's say f(x) = x^2

OpenStudy (anonymous):

Then the slope of the line between (0, f(0) and (.5, f(.5)) is (.25-0)/(.5-0) = 0.5

OpenStudy (anonymous):

next find where the derivative equals 0.5. This is the point where the tangent line is parallel to the secant line between (0, f(0) and (.5, f(.5))

OpenStudy (anonymous):

sorry I crashed too haha

OpenStudy (anonymous):

so f'(x) = 2x = 0.5 x = 0.25

OpenStudy (anonymous):

so is d 0.25?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

well, no actually

OpenStudy (anonymous):

d is 0.5

OpenStudy (anonymous):

oh k, that makes more sense

OpenStudy (anonymous):

because f'(0.25) = 0.5 so c = 0.25 d = 0.5

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

no problem